OpenStudy (anonymous):

Vectors. Let A, B and C be collinear and such that AB : BC = 3 : 2. Find AB : BC expressing your answer in terms of integers

OpenStudy (anonymous):

AB : BC = 3:2 2AB = 3BC AC = AB + BC 2/3BC + BC = 5/2BC = -5/2CB Then I don't know what to do :S

OpenStudy (anonymous):

what is : ?

OpenStudy (anonymous):

nm I think I worked it out 2AB = -5CB so AC : CB = -5:2